The most recent financial statements for Williamson, Inc., are shown here (assuming no income taxes):

Income Statement Balance Sheet
  Sales $ 9,300   Assets $ 20,000   Debt $ 8,000
  Costs 7,330   Equity 12,000
    Net income $ 1,970     Total $ 20,000     Total $ 20,000

Assets and costs are proportional to sales. Debt and equity are not. No dividends are paid. Next year’s sales are projected to be $10,788.

What is the external financing needed? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

Growth rate in sales=(10788-9300)/9300=16%

Sales 10788
Costs(7330*1.16) 8502.8
Net income 2285.2

Total assets would be=20,000*1.16=23200

Total equity would be=12000+Net income

=12000+2285.2=14285.2

Total assets=Total equity+Total debt

Hence external financing needed=23200-(14285.2+8000)

=$915(Approx).